OpenStudy (anonymous):

Solve the Triangle: A = 51°, b = 41, c = 40

OpenStudy (anonymous):

Apply Cosines Law, since you have SAS case

OpenStudy (anonymous):

c^2 = a^2 + b^2 – 2ab cos C and a^2 = b^2 + c^2 – 2bc cos A and b^2 = a^2 + c^2 – 2ac cos B. Which one would I use?

OpenStudy (anonymous):

Given b, and c => find a

OpenStudy (anonymous):

a² = b² + c² – 2bc cos A

OpenStudy (anonymous):

cos A = (b^2+c^2-a^2)/2bc

OpenStudy (anonymous):

No, this time you find the third side a!

OpenStudy (anonymous):

41^2+40^2=a^2 3281=a^2 sq rt (3281)=57.3 a= 57

OpenStudy (anonymous):

Check your formula again!

OpenStudy (anonymous):

Lol Oh Sorry this time i wouldn't have to change it would I? just plug in numbers?

OpenStudy (anonymous):

a² = b² + c² – 2bc cos A

OpenStudy (anonymous):

a² = b² + c² – 2bc cos A a²=1681+1600-3280*.629 a²=3281-2063 a²=1218 a=34.89

OpenStudy (anonymous):

Correct, I guess you're not familiar with using the calculator, so it takes you so long to get the result !

OpenStudy (anonymous):

no i dont know how to use this calculator lol its new and there is alot more buttons then my old one. haha. its hard to find some things on this one.

OpenStudy (anonymous):

Now, do same thing as your previous post! Apply Sines Law to find the second angle

OpenStudy (anonymous):

i have ten minutes to get the rest of this one and the other two triangles done. can you give my the answers then we can go back and work threw them? if not thats fine ill just turn in what i have. but if so that would be really awesome.

OpenStudy (anonymous):

I tend to forget, so by the time you answer I already forgot what is the next step :(

OpenStudy (anonymous):

so B=sin b/b?

OpenStudy (anonymous):

The second angle is 65.98

OpenStudy (anonymous):

Then it's so easy to find the rest!

OpenStudy (anonymous):

here is the next triangle

OpenStudy (anonymous):

3. A = 23°, B = 55°, b = 9

OpenStudy (anonymous):

C=102 degrees

OpenStudy (anonymous):

@CynthiaJune94 Please open the new post for each question!

OpenStudy (anonymous):

a=sin A/A right?